Fairfax Paint stock is expected to pay a dividend of 4.56 dollars in 1 year and a dividend of 4.07 dollars in 2 years. The stock is currently priced at 129.64 dollars and is expected to be priced at 135.32 dollars in 1 year. What is the price of Fairfax Paint stock expected to be in 2 years? The stock’s dividend is paid annually and the next dividend is expected in 1 year.
Answer #1

Price of Stock today = (Dividend in Year 1+ Price in Year 1)/(1+r)
129.64 = (4.56 + 135.32)/(1+r)
(1+r) = (4.56+135.32)/129.64
r = 7.8988%

Price of stock next year = (Dividend in Year 2+ Price in Year 2)/(1+r)
135.32 = (4.07 + Price in year 2)/(1+7.8999%)
Price in year 2 = 135.32*(1+7.8999%)-4.07 = 141.94

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well